3) Consider a German firm taht wishes to invest euro funds for a period of one year. the firm has a choice of investing in a euro bond with one year to maturity, paying an interest rate of 3.35 percent and U.S. dollar bond with one year to maturity, paying an interest rate of 2.25 percent. The current exchange rate i:s euro 1.25 per dollar. Should the german firm invest in euro bond or in the US bond? Explain
Answer #1
Given,the current exchange rate of Euro to US Dollar : 1 Euro will fetch 1.25 US $
So, an amount of 1000 Euros is equal to 1000*1.25= 1250 US $
Interest amt. on Euro bond on the amount invested=1000*3.35%= 33.5 euros
Interest on US bond on the amount invested=1250*2.25%= 28.13 which is (1/1.25*28.13)= 22.50 euros
33.5>22.5
So, the German firm should invest in Euro bonds only.